OpenStudy (anonymous):

Planet A and Planet B are in circular orbits around a distant star. Planet A is 6.0 times farther from the star than Planet B. Find the ratio of their speeds. Va/Vb.

OpenStudy (mos1635):

aplay centripatal force condition for circular motion with only gravitational force, the each planet do you see how to proceed??

OpenStudy (anonymous):

 hey tell me first the diagram i drawn is correct or not??

OpenStudy (anonymous):

here u ll see the force is gravitational force now we will compare it with cetrepetal force

OpenStudy (anonymous):

let mass of star=m mass of planet A =m1 mass of planet B=m2 F=Gm1m/r^2.......................(1) centrpetal force F=mvA^2/r.............(2) by comparing equation 1 and 2 vA= sqrtGm1/rA similary vB=sqrt Gm2/rB u are given dat rA=6rB by solving u ll get vA/vB=sqrt m1/6m2

OpenStudy (anonymous):

So its \[\sqrt{1/6}= .4082?\]

OpenStudy (anonymous):

but u are not given mass so how can we asume that m1 and m2 are equal so in ratio u have to put m1 and m2 too..

OpenStudy (mos1635):

no no no heena in your comparition of equation 1 and 2 you have M(of planet) on your final formula.

OpenStudy (mos1635):